The electric field vectors of two light waves propagating along the positive zdirection

are given as

E1(z,t) = xˆ E01cos(kz - (omega)t)

E2(z,t) = yˆ E02cos(kz-(omega)t + (si))

where xˆ and yˆ are unit vectors along x and y-axes, respectively. Show that when

these two waves superpose, we obtain elliptically polarised light. Also show that

the linear and circular polarisations are special cases of elliptical polarisation


1
Expert's answer
2021-05-05T14:50:45-0400

We have given electric field vectors of two light waves propagating along the positive z-direction.


"E_1(z,t) = xˆ E_{01}cos(kz - \omega t)-----(1)"

"E_2(z,t) = yˆ E_{02}cos(kz-\omega t + \phi)-----(2)"


Where, E_{01} and E_{02} are amplitudes, \omega is cyclic frequency ,

From equation (1),

"cos(kz- \omega t) = \dfrac{E_1}{E_{01}}-----(3)"

From equation (2),


"E_2 = E_{02}[cos(kz- \omega t)cos \phi - sin(kz- \omega t)sin \phi]-----(4)"


From equation (4),


"sin(kz-\omega t)sin \phi = cos(kz - \omega t)cos\phi - \dfrac{E_2}{E_{02}}-----(5)"


Putting (3) equation in (5), We get:


"sin(kz- \omega t)sin \phi = \dfrac{E_1}{E_{01}} cos\phi - \dfrac{E_2}{E_{02}}----(6)"


Squaring both sides,


"(sin(kz- \omega t)sin \phi)^2 = (\dfrac{E_1}{E_{01}} cos\phi - \dfrac{E_2}{E_{02}})^2---(7)"


"(sin(kz- \omega t)sin \phi)^2 = (\dfrac{E_1cos \phi}{E_{01}})^2+ (\dfrac{E_2}{E_{02}})^2 - 2\times \dfrac{E_1 cos \phi}{E_{01}} \times \dfrac{E_2}{E_{02}}---(8)"


From equation (3),


Multiplying both sides by "sin \phi" we get,

"cos(kz- \omega t)sin\phi = \dfrac{E_1}{E_{01}} sin\phi---(9)"

Squaring both sides,


"(cos(kz- \omega t)sin\phi)^2 = (\dfrac{E_1}{E_{01}} sin\phi)^2---(10)"


Adding Equation (8) and (10)


We get,


"sin^2 \phi = (\dfrac{E_1}{E_{01}})^2 + (\dfrac{E_2}{E_{02}})^2 - 2\times \dfrac{E_1 cos \phi}{E_{01}} \times \dfrac{E_2}{E_{02}}"


which is the equation of Ellipse.


In general, projections of light vectors of polarized light on the axes X and Y


"E_x = E_0cos(\omega t- kz)"


"E_y = E_0cos(\omega t - kz+ \phi)"


they satisfy the equation


"(\dfrac{E_x}{E_0})^2 - 2\dfrac{E_xE_y}{E_0^2}cos\phi + (\dfrac{E_y}{E_0})^2 = sin^2 \phi"


This equation is an equation of an ellipse whose axes are oriented relative to the coordinate

axes X and Y arbitrarily. The orientation of the ellipse and magnitude of its semiaxes depends only on the angle "\phi" (phase difference).


If "\phi = \pi"


"E_x + E_y = 0"


we have the equation of the line. That is, light is linearly or plane polarized.


If "\phi = 0"


"E_x^2+ E_y^2 = E_0^2"


We have the equation of the circle. That is, light is circularly polarized.